Mackinaw Inc. processes a base chemical into plastic. Standard costs and actual costs for direct materials, direct labor, and factory overhead incurred for the manufacture of 4,400 units of product were as follows:

Standard Costs Actual Costs
Direct materials 5,700 lb. at $5.50 5,600 lb. at $5.40
Direct labor 1,100 hrs. at $16.60 1,130 hrs. at $16.90
Factory overhead Rates per direct labor hr.,
based on 100% of normal
capacity of 1,150 direct
labor hrs.:
Variable cost, $3.60 $3,920 variable cost
Fixed cost, $5.70 $6,555 fixed cost

Each unit requires 0.25 hour of direct labor.

Required:

a. Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct materials price variance $
Direct materials quantity variance
Total direct materials cost variance $

b. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct labor rate variance $
Direct labor time variance
Total direct labor cost variance $

c. Determine variable factory overhead controllable variance, the fixed factory overhead volume variance, and total factory overhead cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Variable factory overhead controllable variance $
Fixed factory overhead volume variance
Total factory overhead cost variance $
Answer #1

a) Direct Materials Price Variance = (Std Price - Actual Price)*Actual Qty

= ($5.50 - $5.40)*5,600 lbs. = $560 Favorable

Direct Materials Quantity Variance = (Std Qty - Actual Qty)*Std Price

= (5,700 - 5,600)*$5.50 = $550 Favorable

Total Direct materials cost variance = $560 F + $550 F = $1,110 Favorable

b) Direct Labor Rate Variance = (Std Rate - Actual Rate)*Actual Hrs

= ($16.60 - $16.90)*1,130 hrs = $339 Unfavorable

Direct Labor Time Variance = (Std Hrs - Actual Hrs)*Std Rate

= (1,100 - 1,130)*$16.60 = $498 Unfavorable

Total Direct labor cost variance = $339 U + $498 U = $837 Unfavorable

c) Variable overhead spending variance = Actual Variable Overhead - (Actual hrs*Std variable OH rate)

= $3,920 - (1,130 hrs*$3.60) = $148 Favorable

Variable overhead efficiency variance = (Std hrs - Actual hrs)*Std variable overhead rate

= (1,100 - 1,130)*$3.60 = $108 Unfavorable

Variable overhead controllable variance = $148 F+$108 U = $40 F

Budgeted Fixed Overheads = 1,150 hrs*$5.70 = $6,555

Fixed Overhead Budget Variance = Actual Fixed Overhead - Budgeted Fixed Overheads

= $6,555 - $6,555 = $0

Fixed Overhead Volume Variance = Budgeted Fixed Overheads - Applied Fixed Overheads

= $6,555 - (1,130*$5.70) = $114 U

Variable factory overhead controllable variance $40 F
Fixed factory overhead volume variance $114 U
Total factory overhead cost variance $74 U